what is the magnitude of the momemtum of a 1000kg car whose speed is 20.0m/s

Grade:upto college level

1 Answers

Sher Mohammad IIT Delhi
9 years ago
momentum(P)=mass*velocity
so P=20*1000 kg m/sec=2*10^4 newton-sec

Sher mohammad
faculty askiitian
B.tech IIT delhi